Four coins are tossed. Let Z be the random variable representing the number of heads that occur. Find the values of the random variable Z. Construct a probability distribution and draw the corresponding histogram


1
Expert's answer
2022-03-08T07:18:04-0500

Z ~ BIN(4, 0.5), so Z = {0, 1, 2, 3, 4}

"P(Z=0)={4 \choose 0}*({\frac 1 2})^4={\frac 1 {16}}"

"P(Z=1)={4 \choose 1}*({\frac 1 2})^4={\frac 4 {16}}={\frac 1 4}"

"P(Z=2)={4 \choose 2}*({\frac 1 2})^4={\frac 6 {16}}={\frac 3 8}"

"P(Z=3)={4 \choose 3}*({\frac 1 2})^4={\frac 4 {16}}={\frac 1 4}"

"P(Z=4)={4 \choose 4}*({\frac 1 2})^4={\frac 1 {16}}"